# Coarse-to-fine Q-Network with Action Sequence for Data-Efficient Robot Learning

### v3

@@ Abstract
In reinforcement learning (RL), we train a value function to understand the long-term consequence of executing a single action. However, the value of taking each action can be ambiguous in robotics as robot movements are typically the aggregate result of executing multiple small actions. Moreover, robotic training data often consists of noisy trajectories, in which each action is noisy but executing a series of actions results in a meaningful robot movement. This further makes it difficult for the value function to understand the effect of individual actions. To address this, we introduce Coarse-to-fine Q-Network with Action Sequence (CQN-AS), a novel value-based RL algorithm that learns a critic network that outputs Q-values over a sequence of actions, i.e., explicitly training the value function to learn the consequence of executing action sequences. We study our algorithm on 53 robotic tasks with sparse and dense rewards, as well as with and without demonstrations, from BiGym, HumanoidBench, and RLBench. We find that CQN-AS outperforms various baselines, in particular on humanoid control tasks.
Project webpage: younggyo.me/cqn-as
Pieter Abbeel
University of California, Berkeley
mail@younggyo.me
@@ 1 Introduction
Reinforcement learning (RL) holds the promise of continually improving policies through online experiences (Sutton & Barto, 2018). However, training RL agents on robotic tasks often requires a prohibitively large number of training samples (Kalashnikov et al., 2018; Herzog et al., 2023), which is problematic as deploying robots incurs a huge cost.
We posit that one cause for the poor data-efficiency of RL for robotics is the credit assignment problem (Minsky, 1961). In value-based RL, we typically train a value function to understand the long-term consequences of taking individual actions. However, this becomes challenging in robotics as robot movements are the aggregate result of a sequence of low-level actions, where the effect of each action is often ambiguous. For instance, moving a robot arm forward involves multiple steps of changing joint angles. While one may easily understand the consequence of executing a series of actions, it is often difficult to learn how each low-level action, such as changing the joint angle by certain degrees, contributes to the resulting movement.
Figure 1: Noisy trajectories in robotic data. We visualize the $(x,y,z)$ coordinates of a gripper in demonstrations from RLBench (James et al., 2020) that uses motion-planning for generating demonstrations and BiGym (Chernyadev et al., 2024) that provides human-collected demonstrations. (Right) We report the success rates of replaying BiGym demonstrations with various action smoothing schemes. We find that naïve approach of smoothing actions can make demonstrations invalid as smoothed actions often lose precision, highlighting the need for developing RL algorithms that can learn from noisy robotic training data.
Moreover, robotic training data often consists of noisy trajectories, amplifying the credit assignment problem. For instance, we often initialize training with human-collected demonstrations that consist of noisy multi-modal trajectories (Chernyadev et al., 2024). In such trajectories, each action may be noisy because of human errors; however, executing a sequence of actions results in meaningful robot movement. One may think that this issue can be resolved by smoothing actions, but it often makes demonstrations invalid by reducing action precision (see Figure 1). Moreover, when training RL agents, we typically inject some noise into actions for exploration (Sehnke et al., 2010; Lillicrap et al., 2016), which induces trajectories with jerky motions. This characteristic of robotic training data further makes it difficult to learn the value function that can properly evaluate the long-term consequence of each individual action.
@@
On the other hand, recent behavior cloning (BC; Pomerleau 1988) approaches have shown that predicting a sequence of actions enables policies to effectively approximate the noisy, multi-modal distribution of expert demonstrations (Zhao et al., 2023; Chi et al., 2023). Inspired by this, we investigate whether we can similarly train a value function to explicitly learn the consequence of executing action sequences.
@@
As a straightforward implementation of this idea, in Figure 2, we train actor-critic algorithms (Haarnoja et al., 2018; Fujimoto et al., 2018) with action sequence on stand task from HumanoidBench (Sferrazza et al., 2024). Specifically, we train the actor to output action sequence and the critic to take action sequence as inputs instead of single-step actions. Unfortunately, we observe that actor-critic algorithms with action sequence suffer from a severe value overestimation. This is because a wider action space makes the critic more vulnerable to function approximation error (Fujimoto et al., 2018) and the actor can easily exploit the estimation error.
@@
These results motivate us to design our RL algorithm with action sequence upon a recent value-based algorithm, i.e., Coarse-to-fine Q-Network (CQN; Seo et al. 2024), which solves continuous control tasks with discrete actions. We find that training this critic-only algorithm with action sequence enables us to achieve the benefits of RL with action sequence yet avoid the value overestimation problem. In particular, we introduce Coarse-to-fine Q-Network with Action Sequence (CQN-AS), which learns a critic network that outputs Q-values over a sequence of actions (see Figure 3). By training the critic network to explicitly learn the consequence of taking a sequence of current and future actions, CQN-AS enables the RL agents to effectively learn useful value functions on challenging robotic tasks.
We show that CQN-AS outperforms various baselines on diverse setups with sparse and dense rewards, as well as with or without demonstrations: (i) mobile bi-manual manipulation tasks from BiGym (Chernyadev et al., 2024) that provides human-collected demonstrations, (ii) densely-rewarded humanoid control tasks from HumanoidBench (Sferrazza et al., 2024), and (iii) table-top manipulation tasks from RLBench (James et al., 2020) that provides synthetic demonstrations generated via motion-planning.
Figure 2: Training actor-critic algorithms with action sequences leads to value overestimation. We report target Q-values recorded throughout training actor-critic algorithms, i.e., SAC (Haarnoja et al., 2018) and TD3 (Fujimoto et al., 2018), with action sequence on $\tt{stand}$ task from HumanoidBench (Sferrazza et al., 2024).
@@ 2 Preliminaries
(a) Coarse-to-fine inference procedure
(b) Architecture
Figure 3: Coarse-to-fine Q-Network with Action Sequence (CQN-AS). We build our algorithm upon Coarse-to-fine Q-Network (CQN; Seo et al. 2024), a recent critic-only RL algorithm that solves continuous control tasks with discrete actions. (a) In CQN framework, we train RL agents to zoom-into the continuous action space by iterating the procedures of (i) discretizing continuous action space into $B$ bins and (ii) find the bin with the highest Q-value to further discretize at the next level. We then use the last level’s action sequence for controlling robots. CQN-AS extends this idea to action sequences by computing actions for all sequence steps $k\in[1,...,K]$ in parallel. (b) We train a critic network to output Q-values over a sequence of actions. We design our architecture to first obtain features for each sequence step and aggregate features from multiple sequence steps with a recurrent network. We then project these outputs into Q-values.
@@
Coarse-to-fine Q-Network (CQN; Seo et al. 2024) is a value-based RL algorithm that solves continuous control tasks with discrete actions. The main idea of CQN is to train an RL agent to learn to select coarse discrete actions in shallower levels with larger bin sizes, and then refine their choices by selecting finer-grained actions in deeper levels with smaller bin sizes. Specifically, CQN iterates the procedures of (i) discretizing the continuous action space into multiple bins and (ii) selecting the bin with the highest Q-value to further discretize. This reformulates the continuous control problem as a multi-level discrete control problem, allowing for the use of ideas from sample-efficient discrete RL algorithms (Mnih et al., 2015; Silver et al., 2017) for continuous control.
Formally, let $a_{t}^{l}$ be an action at level $l$ with $a_{t}^{0}$ being the zero vector. [^1] For simplicity, we describe CQN and CQN-AS with a single-dimensional action in the main section. See Appendix B for full description with $N$-dimensional actions. We then define the coarse-to-fine critic to consist of multiple Q-networks which compute Q-values for actions at each level $a_{t}^{l}$, given the features $\mathbf{h}_{t}$ and actions from the previous level $a_{t}^{l-1}$, as follows:
$Q^{l}_{\theta}(\mathbf{h}_{t},a^{l-1}_{t})=\left[Q^{l}_{\theta}(\mathbf{h}_{t},a^{l}_{t}=a^{l,b}_{t},a^{l-1}_{t})\right]_{b=1}^{B}\in\mathbb{R}^{B}$
@@
We then use the last level’s action $a_{t}^{L}$ as the action at time step $t$. For more details, including the inference procedure for computing Q-values, we refer readers to Appendix B.
@@ 3 Method
Figure 4: Examples of robotic tasks. We study CQN-AS on 53 tasks spanning mobile bi-manual manipulation, whole-body control, and tabletop manipulation from BiGym (Chernyadev et al., 2024), HumanoidBench (Sferrazza et al., 2024), and RLBench (James et al., 2020).
@@ 3.1 Coarse-to-fine Critic with Action Sequence
where $N$ is a hyperparameter for $N$-step return and $\pi_{K}^{l}$ is an action sequence policy that outputs the action sequence $\mathbf{a}_{t:t+K}^{l}$ by following the similar inference procedure as in Section 2 (see Figure 3(a)). In practice, we compute Q-values for all sequence step $k\in\{1,...,K\}$ in parallel, which is possible as Q-values for future actions depend only on features $\mathbf{h}_{t}$ but not on previous actions.
@@
We note that any $N$-step return can be used in Equation 1 because the network can learn the long-term value of outputting action $a_{t+k}$ from bootstrapping. There is a trade-off: if one considers a short $N$-step return, it can cause a challenge as the setup becomes a delayed reward setup; but training with higher $N$-step return may introduce variance (Sutton & Barto, 2018). In our considered setups, we empirically find that using common values $N\in\{1,3\}$ works the best. We provide empirical analysis on the effect of $N$ in Figure 8(c).
@@
Our critic network initially extracts features for each sequence step $k$ and aggregates features from multiple steps with a recurrent network (see Figure 3(b)). This architecture is helpful in cases where a single-step action is already high-dimensional so that concatenating them make inputs too high-dimensional. Specifically, let $\mathbf{e}_{k}$ denote an one-hot encoding for $k$. At each level $l$, we construct features for each sequence step $k$ as $\mathbf{h}^{l}_{t,k}=\left[\mathbf{h}_{t},a_{t+k-1}^{l-1},\mathbf{e}_{k}\right]$. We then encode each $\mathbf{h}_{t,k}^{l}$ with a shared MLP network and process them through GRU (Cho et al., 2014) to obtain $\mathbf{s}^{l}_{t,k}=f^{\tt{GRU}}_{\theta}(f^{\tt{MLP}}_{\theta}(\mathbf{h}_{t,1}^{l}),...,f^{\tt{MLP}}_{\theta}(\mathbf{h}_{t,k}^{l})$. We find that this design empirically performs better than directly giving actions as inputs to GRU. We then use a shared projection layer to map each $\mathbf{s}_{t,k}^{l}$ into Q-values at each sequence step $k$, i.e., $[Q_{\theta}^{l,k}(\mathbf{h}_{t},a_{t+k-1}^{l,b},a_{t:t+K}^{l-1})]_{b=1}^{B}=f_{\theta}^{\tt{proj}}(\mathbf{s}_{t,k}^{l})$.
@@ 3.2 Action Execution and Training Details
Figure 5: BiGym results on 25 sparsely-rewarded mobile bi-manual manipulation tasks. All experiments are initialized with 17 to 60 human-collected demonstrations, and RL methods are trained with an auxiliary BC objective. On many tasks, CQN-AS quickly matches the performance of ACT (Zhao et al., 2023) and surpasses it through online learning, also outperforming other RL baselines. We report the success rate over 25 episodes. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 8 runs.
@@ 4 Experiment
We study CQN-AS on 53 robotic tasks spanning mobile bi-manual manipulation, whole-body control, and tabletop manipulation tasks from BiGym (Chernyadev et al., 2024), HumanoidBench (Sferrazza et al., 2024), and RLBench (James et al., 2020) environments (see Figure 4 for examples of robotic tasks). These tasks with sparse and dense rewards, with or without vision sensors, and with or without demonstrations, allow for evaluating the capabilities and limitations of our algorithm. In particular, our experiments are designed to investigate the following questions:
@@
    Can CQN-AS quickly match the performance of a recent BC algorithm (Zhao et al., 2023) and surpass it through online learning? How does CQN-AS compare to previous model-free RL algorithms (Haarnoja et al., 2018; Yarats et al., 2022; Seo et al., 2024)?
@@
Baselines for tasks with demonstrations
For manipulation tasks from BiGym and RLBench, we consider model-free RL baselines that learn deterministic policies, as we find that stochastic policies struggle to solve such fine-grained control tasks. Specifically, we consider (i) Coarse-to-fine Q-Network (CQN; Seo et al. 2024), our backbone algorithm and (ii) DrQ-v2+, an optimized demo-driven variant of an actor-critic algorithm DrQ-v2 (Yarats et al., 2022) that uses a deterministic policy algorithm and data augmentation. We further consider (iii) Action Chunking Transformer (ACT; Zhao et al. 2023), a BC algorithm that trains a transformer (Vaswani et al., 2017) policy to predict action sequence and utilizes temporal ensemble, as our BC baseline.
Baselines for humanoid control tasks with dense reward
For HumanoidBench tasks, we consider (i) Soft Actor-Critic (SAC; Haarnoja et al. 2018), a model-free actor-critic RL algorithm that maximizes action entropy and (ii) CQN.
@@
Figure 6: HumanoidBench results on 8 densely-rewarded humanoid control tasks (Sferrazza et al., 2024). All the experiments start from scratch and RL methods do not have an auxiliary BC objective. CQN-AS significantly outperforms other model-free RL baselines on most tasks. For CQN-AS and CQN, we report the results aggregated over 8 runs. For SAC, we report the results aggregated over 3 runs available from public website. The solid line and shaded regions represent the mean and confidence intervals.

### v4

@@ Abstract
Predicting a sequence of actions has been crucial in the success of recent behavior cloning algorithms in robotics. Can similar ideas improve reinforcement learning (RL)? We answer affirmatively by observing that incorporating action sequences when predicting ground-truth return-to-go leads to lower validation loss. Motivated by this, we introduce Coarse-to-fine Q-Network with Action Sequence (CQN-AS), a novel value-based RL algorithm that learns a critic network that outputs Q-values over a sequence of actions, i.e., explicitly training the value function to learn the consequence of executing action sequences. Our experiments show that CQN-AS outperforms several baselines on a variety of sparse-reward humanoid control and tabletop manipulation tasks from BiGym and RLBench.
Figure 1: Summary of results. Coarse-to-fine Q-Network with Action Sequence (CQN-AS) learns a critic network with action sequence. CQN-AS outperforms various RL and BC baselines such as CQN (Seo et al., 2024), DrQ-v2+ (Yarats et al., 2022), and ACT (Zhao et al., 2023) on 45 robotic tasks from BiGym (Chernyadev et al., 2024) and RLBench (James et al., 2020).
@@ 1 Introduction
Predicting action sequences from expert trajectories is a key idea in recent successful behavior cloning (BC; Pomerleau 1988) approaches in robotics. This has enabled policies to effectively approximate the noisy, multi-modal distribution of expert demonstrations (Zhao et al., 2023; Chi et al., 2023). Can this idea similarly be useful for reinforcement learning (RL)?
Our initial finding is affirmative: we make an intriguing observation that using action sequences can enhance value learning. Specifically, with humanoid demonstrations from BiGym (Chernyadev et al., 2024), we train regression models that predict the ground-truth return-to-go, i.e., the sum of discounted future rewards from the timestep $t$, given the current observation and action. In Figure 2a, we find that using an action sequence $\mathbf{a}_{t:t+K}=\{\mathbf{a}_{t},...,\mathbf{a}_{t+K-1}\}$ as input results in lower validation losses than using a single-step action $\mathbf{a}_{t}$. We hypothesize this is because action sequences, which can correspond to behavioral primitives such as going straight, make it easier for the model to learn the long-term outcomes compared to evaluating the effect of individual single-step actions.
@@
Building on this observation, we train actor-critic algorithms (Haarnoja et al., 2018; Fujimoto et al., 2018) with action sequence on stand task from HumanoidBench (Sferrazza et al., 2024). Specifically, we train the actor to output action sequence and the critic to take action sequence as inputs instead of single-step actions. However, we find these algorithms with action sequences suffer from severe value overestimation (see Figure 2b) and completely fail to solve the task (see Figure 2c). This is because a wider action space makes the critic more vulnerable to function approximation error (Fujimoto et al., 2018) and the actor excessively maximizes value functions by exploiting this estimation error.
(a) Return-to-go prediction
(b) Value overestimation on stand
(c) Performance on stand
Figure 2: Analyses. (a) We measure the improvement in the validation L1 loss of the return-to-go prediction model with different action sequence lengths. We find that using action sequence of length 50 results in the lower loss than using single-step action. (b) We find that SAC and TD3 with action sequences suffer from severe value overestimation in stand task from HumanoidBench (Sferrazza et al., 2024). (c) This leads to unstable training and failure to solve the task. These findings motivate us to design CQN-AS, a novel RL algorithm that incorporates action sequences for value learning while avoiding the value overestimation problem.
@@
This result motivates us to design our RL algorithm with action sequence upon a recent critic-only algorithm, i.e., Coarse-to-fine Q-Network (CQN; Seo et al. 2024), which solves continuous control tasks with discrete actions. Because there is no separate actor that may exploit value functions, i.e., CQN simply selects discrete actions with the highest Q-values, we find that training with action sequences is stable and thus avoids value overestimation problem. In particular, we introduce Coarse-to-fine Q-Network with Action Sequence (CQN-AS), which learns a critic network that outputs Q-values over a sequence of actions (see Figure 3). By training the critic network to explicitly learn the consequence of taking a sequence of current and future actions, CQN-AS enables the RL agents to effectively learn useful value functions on challenging robotic tasks.
Our experiments show that CQN-AS improves the performance of CQN on sparse-reward humanoid control tasks from BiGym benchmark (Chernyadev et al., 2024) that provides human-collected demonstrations and sparse-reward tabletop manipulation tasks from RLBench (James et al., 2020) that provide demonstrations generated via motion-planning. Considering that CQN-AS is a critic-only algorithm that selects actions with the highest Q-value without a separate actor network, these results highlight the benefit of using action sequences in value learning.
Our contributions can be summarized as below:
  • We make an observation that shows using action sequences can be useful for RL by enhancing value learning. We also show that standard actor-critic algorithms (Haarnoja et al., 2018; Fujimoto et al., 2018) suffer from value overestimation when trained with action sequences.
  • We introduce Coarse-to-fine Q-Network with Action Sequence (CQN-AS) that trains a critic network to output Q-values over action sequences. This critic-only algorithm successfully avoids value overestimation problem and enhances the base CQN algorithm.
  • In a demo-driven RL setup that initializes training with expert demonstrations, we show that CQN-AS surpasses the performance of ACT (Zhao et al., 2023) – a BC algorithm that trains a Transformer (Vaswani et al., 2017) to predict action sequences.
@@ 2 Preliminaries
(a) Coarse-to-fine inference procedure
(b) Architecture
Figure 3: Coarse-to-Fine Q-Network with Action Sequence. CQN-AS extends Coarse-to-Fine Q-Network (CQN; Seo et al. 2024), a critic-only RL algorithm for continuous control using discretized actions. (a) CQN progressively zooms into the action space by discretizing it into $B$ bins and finding the bin with the highest Q-value to further discretize at the next level. Last level’s action sequence is used for controlling robots. CQN-AS generalizes this to action sequences by computing all $K$ actions in parallel. (b) We train a critic to predict Q-values over entire action sequences by extracting per-step features and aggregating them with a recurrent network before projection to Q-values.
@@
Coarse-to-fine Q-Network (CQN; Seo et al. 2024) is a critic-only RL algorithm that solves continuous control tasks with discrete actions. CQN trains an RL agent to learn to select coarse discrete actions in shallower levels with larger bin sizes, and then refine their choices by selecting finer-grained actions in deeper levels with smaller bin sizes. Specifically, CQN iterates the procedures of (i) discretizing the continuous action space into multiple bins and (ii) selecting the bin with the highest Q-value to further discretize. This reformulates the continuous control problem as a multi-level discrete control problem, allowing for the use of ideas from sample-efficient discrete RL algorithms (Mnih et al., 2015; Silver et al., 2017) for continuous control.
Formally, let $a_{t}^{l}$ be an action at level $l$ with $a_{t}^{0}$ being the zero vector. [^1] For simplicity, we describe CQN and CQN-AS with a single-dimensional action in the main section. See Appendix C for full description with $N$-dimensional actions. We then define the coarse-to-fine critic to consist of multiple Q-networks which compute Q-values for actions at each level $a_{t}^{l}$, given the features $\mathbf{h}_{t}$ and actions from the previous level $a_{t}^{l-1}$, as follows:
$Q^{l}_{\theta}(\mathbf{h}_{t},a^{l-1}_{t})=\left[Q^{l}_{\theta}(\mathbf{h}_{t},a^{l}_{t}=a^{l,b}_{t},a^{l-1}_{t})\right]_{b=1}^{B}\in\mathbb{R}^{B}$ (1)
@@
We then use the last level’s action $a_{t}^{L}$ as the action at time step $t$. For more details, including the inference procedure for computing Q-values, we refer readers to Appendix C.
@@ 3.1 Coarse-to-fine Critic with Action Sequence
where $N$ is a hyperparameter for $N$-step return and $\pi_{K}^{l}$ is an action sequence policy that outputs the action sequence $\mathbf{a}_{t:t+K}^{l}$ by following the similar inference procedure as in Section 2 (see Figure 3a). In practice, we compute Q-values for all sequence step $k\in\{1,...,K\}$ in parallel, which is possible as Q-values for future actions depend only on features $\mathbf{h}_{t}$ but not on previous actions.
@@
We note that any $N$-step return can be used in Equation 2 because the network can learn the long-term value of outputting action $a_{t+k}$ from bootstrapping. There is a trade-off: if one considers a short $N$-step return, it can cause a challenge as the setup becomes a delayed reward setup; but training with higher $N$-step return may introduce variance (Sutton & Barto, 2018). In our considered setups, we empirically find that using common values $N\in\{1,3\}$ works the best. We provide empirical analysis on the effect of $N$ in Figure 7c.
@@
Our critic network initially extracts features for each sequence step $k$ and aggregates features from multiple steps with a recurrent network (see Figure 3b). This architecture is helpful in cases where a single-step action is already high-dimensional so that concatenating them make inputs too high-dimensional. Specifically, let $\mathbf{e}_{k}$ denote an one-hot encoding for $k$. At each level $l$, we construct features for each sequence step $k$ as $\mathbf{h}^{l}_{t,k}=\left[\mathbf{h}_{t},a_{t+k-1}^{l-1},\mathbf{e}_{k}\right]$. We then encode each $\mathbf{h}_{t,k}^{l}$ with a shared MLP network and process them through GRU (Cho et al., 2014) to obtain $\mathbf{s}^{l}_{t,k}=f^{\tt{GRU}}_{\theta}(f^{\tt{MLP}}_{\theta}(\mathbf{h}_{t,1}^{l}),...,f^{\tt{MLP}}_{\theta}(\mathbf{h}_{t,k}^{l})$. We find that this design empirically performs better than directly giving actions as inputs to GRU. We then use a shared projection layer to map each $\mathbf{s}_{t,k}^{l}$ into Q-values at each sequence step $k$, i.e., $[Q_{\theta}^{l,k}(\mathbf{h}_{t},a_{t+k-1}^{l,b},a_{t:t+K}^{l-1})]_{b=1}^{B}=f_{\theta}^{\tt{proj}}(\mathbf{s}_{t,k}^{l})$.
@@ 3.2 Action Execution and Training Details
Figure 4: Examples of robotic tasks. We study CQN-AS on 25 humanoid control tasks from BiGym (Chernyadev et al., 2024) and 20 tabletop manipulation tasks from RLBench (James et al., 2020).
@@ 4 Experiment
We study CQN-AS on 25 humanoid control tasks from BiGym (Chernyadev et al., 2024) and 20 tabletop manipulation tasks from RLBench (James et al., 2020) (see Figure 4 for examples of robotic tasks). To focus on challenging robotic tasks that aim to induce policies generating realistic behaviors, we consider a practical setup of demo-driven RL where we initialize training with a modest amount of expert demonstrations and then train with online data. In particular, our experiments are designed to investigate the following questions:
@@
    Can CQN-AS quickly match the performance of a recent BC algorithm (Zhao et al., 2023) and surpass it through online learning? How does CQN-AS compare to previous model-free RL algorithms (Yarats et al., 2022; Seo et al., 2024)?
@@
Baselines
We consider model-free RL baselines that learn deterministic policies, as we find that stochastic policies struggle to solve fine-grained manipulation tasks. Specifically, we consider (i) Coarse-to-fine Q-Network (CQN; Seo et al. 2024), our backbone algorithm and (ii) DrQ-v2+, an optimized demo-driven variant of an actor-critic algorithm DrQ-v2 (Yarats et al., 2022) that uses a deterministic policy algorithm and data augmentation. We further consider (iii) Action Chunking Transformer (ACT; Zhao et al. 2023) that trains a transformer (Vaswani et al., 2017) policy to predict action sequences from expert demonstrations and utilizes temporal ensemble, as our BC baseline.
@@
Figure 5: BiGym results on 25 sparsely-rewarded mobile bi-manual manipulation tasks. All experiments are initialized with 17 to 60 human-collected demonstrations, and RL methods are trained with an auxiliary BC objective. We report the success rate over 25 episodes. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 8 runs.